OpenStudy (anonymous):

2(x-4)^2+2y^2=8 please help me find radius and center

OpenStudy (anonymous):

divide all of it by 2 first

OpenStudy (anonymous):

both sides?

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

how do you divide 2 by (x-4)^2?

OpenStudy (anonymous):

factor 2 out of (x-4)^2 and y^2 so you can divide by 2 on both sides

OpenStudy (anonymous):

oh so dived 2 by x^2-8x+16?

OpenStudy (anonymous):

dont expand (x-4)^2, leave it alone

OpenStudy (anonymous):

would x^2 , the 1st term still be x^2? BECAUSE I CAN DIVIDE THE REST BUT THAT 1ST TERM

OpenStudy (anonymous):

sorry for caps

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

so would it look like : (x-4)^2+2y^2/2=8 /2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now divide it out

OpenStudy (anonymous):

can you give me an example? please , on how to divide (x-4)^2 by 2? byt giving me an example

OpenStudy (anonymous):

you already have in the problem its 2(x-4)^2

OpenStudy (anonymous):

so it would look like 2(x-4)^2+y^2=4?

OpenStudy (anonymous):

it would look like (x-4)^2 + y^2 = 4

OpenStudy (anonymous):

ohhh ok thanks the can you help me get it to standard form and find center and radius and intercepts?

OpenStudy (anonymous):

you can get radius + center from this

OpenStudy (anonymous):

wait , so the center would be (4,1) and radius is 4?

OpenStudy (anonymous):

radius is square root of number on right center is (4, 0) because y^2 is by itself, it doesnt look like (y - 1)^2

OpenStudy (anonymous):

so y^2 is not same as 1y^2? its 0y^2?

OpenStudy (anonymous):

also I got the intercepts as (-3,0)(0,-8)

OpenStudy (lgbasallote):

y^2 is equal to 1y^2 @bunnybee but you have (y-0)^2 so the second coordinate is 0 not y...remember that in looking for the coordinates of the center..you look at the second term of the binomial not the coefficient of the variable ^_^ make any sense?

OpenStudy (anonymous):

yes thank you!